# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 1 Introduction
A system of coalescing random walks (CRW) consists of particles performing i.i.d. rate 1 continuous-time random walks, with immediate coalescence when particles meet. This fundamental model has attracted sustained interest across disciplines due to its dual role as both a mathematically tractable system and a prototypical example connecting diverse physical phenomena. Its duality with the voter model through graphical constructions [Lig05], relevance to chemical reaction kinetics [SFH01], and application to domain wall dynamics in Glauber spin systems [FDM98] underscore its broad significance.
@@
When formulated on $\mathbb{Z}^{d}$, the CRW can be represented as a $\{0,1\}^{\mathbb{Z}^{d}}$-valued Markov process, where occupied (resp. vacant) sites correspond to 1 (resp. 0) states. Consider a system of CRW starting from every site on ${\mathbb{Z}}^{d}$. A central quantity of interest is the particle density $\rho_{1}(t)$, defined via translation invariance as the origin’s occupation probability. Pioneering work by Bramson and Griffeath [BG80], based on a moment calculation, established dimension-dependent asymptotics for simple random walks:
$\rho_{1}(t)\sim\begin{cases}\displaystyle 1/\sqrt{\pi t}&d=1,\\ \displaystyle\log t/(\pi t)&d=2,\\ \displaystyle 1/(\varsigma_{d}t)&d\geq 3,\end{cases}$ (1.1)
@@
where $\varsigma_{d}$ denotes the escape probability for $d$-dimensional simple random walks. Cox and Perkins [CP04] later provided a probabilistic proof using the super-Brownian invariance principle for the dual voter model. For $d\geq 3$, where the simple random walk is transient, van den Berg and Kesten [vdBK00, vdBK02] developed a robust approach by deriving an ordinary differential equation for $\rho_{1}(t)$.
@@
A natural extension of this study is the $N$-point correlation function $\rho_{N}({\bf x},t)$, defined as the occupation probability at sites ${\bf x}:=(x_{1},\ldots,x_{N})$ at time $t$, see (1.8). In the case of coalescing simple random walks in 2D, Lukins, Tribe and Zaboronski [LTZ18] generalized the method of [vdBK00] and proved
$\rho_{N}({\bf x},t)\sim\rho^{N}_{1}(t)\times p_{\rm NC}({\bf x},t),$ (1.2)
@@
where the correction term $p_{\rm NC}({\bf x},t)\sim c_{\bf x}(\log t)^{-n(n-1)/2}$ [CDP00] denotes the non-collision probability of $N$ independent walks starting from ${\bf x}$ by time $t$. This method can be extended to $d\geq 3$, but fails for $d=1$ due to the strong correlations induced by the recurrence of the underlying random walk. Nevertheless, in the case of 1D coalescing Brownian motions, [MRTZ06] derived the asymptotics for $\rho_{N}({\bf x},t)$, relying crucially on the integrability of the system–most notably, the explicit Karlin-McGregor formula.
@@
In contrast to the nearest-neighbor case, much less is known about the $N$-point function for CRW with long-range jumps. The only existing result in this direction is for 1D coalescing $\alpha$-stable processes with $\alpha>1$, where the density $\rho_{1}(t)$ has asymptotic order $t^{1/\alpha}$, derived as a corollary of [EMS13, Thm 6.3].
@@
Motivated by this gap, we study coalescing heavy-tailed random walks on ${\mathbb{Z}}$, where the jump increments lie in the domain of attraction of $\alpha$-stable distributions for $0<\alpha\leq 1$. Our results reveal that a form of (1.2) remains valid in this setting. Notably, $\alpha=1$ is the critical point for recurrence/transience, and our study encompasses a class of recurrent random walks.
@@
Notation. Throughout the paper, we use $c,C$ to denote various positive constants that may even vary within the same line. For two functions $f(t)$ and $g(t)$, we adopt the convention that
@@ 1.1 Model and main result
It is well known [Lig05] that CRW can be constructed by sequentially adding independent random walks, halting each upon its first collision with any of the previously added paths. Precisely, the coalescing heavy-tailed random walks $(\xi_{t})_{t\geq 0}$ on ${\mathbb{Z}}$ can be defined as a $\{0,1\}^{\mathbb{Z}}$-valued Markov process by specifying its generator. For $\xi\in\{0,1\}^{\mathbb{Z}}$ and ${\bf x}=(x_{1},\ldots,x_{n})\in{\mathbb{Z}}^{n}$, write $\xi(x_{1},\ldots,x_{n}):=(\xi(x_{1}),\ldots,\xi(x_{n}))\in\{0,1\}^{n}$. We simplify notation by writing $\xi({\bf x})=1$ if $\xi({\bf x})=(1,\ldots,1)$, and $\xi=1$ if $\xi(x)=1$ for all $x\in{\mathbb{Z}}$. Additionally, we represent elements of $\{0,1\}^{n}$ as binary strings, for example, $(0,1)$ is written as $01$. With this notation, the generator ${\cal G}$ of $(\xi_{t})_{t\geq 0}$ applied on local cylinder functions $f$, i.e., functions that only depend on finitely many coordinates of $\xi$, is given by
$\begin{array}[]{l}\displaystyle{\cal G}f(\xi)=\sum_{x,y\in{\mathbb{Z}}}p(y-x)1_{\{\xi(x,y)=10\}}\big{\{}f(\xi+e_{y}-e_{x})-f(\xi)\big{\}}\\[5.0pt] \hskip 62.59596pt\displaystyle+\sum_{x,y\in{\mathbb{Z}}}p(y-x)1_{\{\xi(x,y)=11\}}\big{\{}f(\xi-e_{x})-f(\xi)\big{\}},\end{array}$ (1.3)
where $e_{x}(y)=1_{\{x=y\}}$, and $p(\cdot)$ with $p(0)=0$ and $\sum_{x}p(x)=1$ denotes the jump kernel, which is in the domain of attraction of an $\alpha$-stable distribution. By [IL71, Thm 2.6.1], it satisfies
$\sum_{|y|\geq x}p(y)=\frac{L(x)}{x^{\alpha}}\big{(}1+o(1)\big{)}\quad\text{as}~{}x\to\infty,$ (1.4)
@@
To avoid unnecessary complexity, we assume that $p$ is symmetric throughout the paper. However, all statements and proofs can be adapted to non-symmetric $p$ without significant difficulty (see Remark 2.8 for further discussion). For the SVF $L$, there exist $1/\alpha$-regularly varying functions $b_{t}:(0,\infty)\to(0,\infty)$ such that $L(b_{t})/b_{t}^{\hskip 1.0pt\alpha}\sim 1/t$. Let $(X_{t})_{t\geq 0}$ be a rate 1 continuous-time random walk with jump kernel $p$ and initial position $X_{0}=0$. It is well known that $X_{t}/b_{t}$ converges weakly to an $\alpha$-stable distribution with probability density function $g_{\alpha}(x)$. By the asymptotic uniqueness of $b_{t}$ [BGT89, Thm 1.5.12] and Gnedenko’s local limit theorem [IL71, Thm 4.2.1], we can choose a specific $1/\alpha$-regularly varying function
$B_{t}:={\mathbb{P}}(X_{t}=0)^{-1}g_{\alpha}(0)$ (1.5)
@@
$\frac{L(B_{t})}{B_{t}^{\alpha}}\sim\frac{1}{t}\quad\mbox{and}\quad\sup_{x\in{\mathbb{Z}}}\big{|}B_{t}{\mathbb{P}}(X_{t}=x)-g_{\alpha}(x/B_{t})\big{|}=o(1).$ (1.6)
Define
@@
Note that $l(t)$ is a SVF for $\alpha\in(0,1]$, while $\ell(t)$ diverges by Karamata’s theorem and is of $(\frac{1}{\alpha}\!-\!1)$-regularly varying. Since $l(\infty)$ and the Green’s function differ only by a constant multiple, an immediate corollary is that the random walk is recurrent if and only if $l(\infty)=\infty$. For $\alpha=1$, we impose an additional condition to control the recurrence behavior.
Function class ${\cal B}$: $B_{t}\in{\cal B}$ if there exists some $K>0$ such that $\ell(t)^{K}\geq l(t)$ for all large $t$.
Obviously, $B_{t}\in{\cal B}$ holds if $l(\infty)<\infty$. A large class of 1-regularly varying functions $B_{t}$ is included in ${\cal B}$. For example, if the SVF $t/B_{t}=l(t)/\ell(t)$ grows slower than $(\log t)^{k}$ for any $k$, then $B_{t}\in{\cal B}$. But if it grows faster than $\exp(\log^{\kappa}t)$ for some $10<\kappa<$ 1, then $B_{t}\notin{\cal B}$.
Our main result concerns the asymptotic behavior of the $N$-point correlation function.
@@
Let $(\xi_{t})_{t\geq 0}$ be the coalescing heavy-tailed random walks on ${\mathbb{Z}}$ with initial configuration $\xi_{0}\equiv 1$. Assume the random walk has jump rate 1 and symmetric jump kernel $p(\cdot)$ of the form (1.4) with $0<\alpha\leq 1$. In the recurrent case with $\alpha=1$ and $l(\infty)=\infty$, assume moreover that $B_{t}\in{\cal B}$. Denote $\upsilon_{N}:={N\choose 2}=\frac{N(N-1)}{2}$. For $N>0$ and ${\bf x}=(x_{1},\ldots,x_{N})\in{\mathbb{Z}}^{N}$, let
$\rho_{N}({\bf x},t):={\mathbb{P}}(\xi_{t}({\bf x})=1)$ (1.8)
@@
be the $N$-point correlation function. Recall $l(t),\ell(t)$ from (1.7), and recall $K$ such that $\ell(t)^{K}\geq l(t)$. Then, there exist $c_{\bf x}$ depending on ${\bf x}$ and some $\eta>0$, such that for any $\varepsilon>0$,

### v2

@@ 1 Introduction
A system of coalescing random walks (CRW) consists of particles performing i.i.d. rate 1 continuous-time random walks, with immediate coalescence when particles meet. This fundamental model has attracted sustained interest across disciplines due to its dual role as both a mathematically tractable system and a prototypical example connecting diverse physical phenomena. Its duality with the voter model through graphical constructions Lig (05), relevance to chemical reaction kinetics SFH (01), and application to domain wall dynamics in Glauber spin systems FDM (98) underscore its broad significance.
@@
When formulated on $\mathbb{Z}^{d}$, the CRW can be represented as a $\{0,1\}^{\mathbb{Z}^{d}}$-valued Markov process, where occupied (resp. vacant) sites correspond to 1 (resp. 0) states. Consider a system of CRW starting from every site on ${\mathbb{Z}}^{d}$. A central quantity of interest is the particle density $\rho_{1}(t)$, defined via translation invariance as the origin’s occupation probability. Pioneering work by Bramson and Griffeath BG (80), based on a moment calculation, established dimension-dependent asymptotics for simple random walks:
$\rho_{1}(t)\sim\begin{cases}\displaystyle 1/\sqrt{\pi t}&d=1,\\ \displaystyle\log t/(\pi t)&d=2,\\ \displaystyle 1/(\varsigma_{d}t)&d\geq 3,\end{cases}$ (1)
@@
where $\varsigma_{d}$ denotes the escape probability for $d$-dimensional simple random walks. Cox and Perkins CP (04) later provided a probabilistic proof using the super-Brownian invariance principle for the dual voter model. For $d\geq 3$, where the simple random walk is transient, van den Berg and Kesten vdBK (00, 02) developed a robust approach by deriving an ordinary differential equation for $\rho_{1}(t)$.
@@
A natural extension of this study is the $N$-point correlation function $\rho_{N}({\bf x},t)$, defined as the occupation probability at sites ${\bf x}:=(x_{1},\ldots,x_{N})$ at time $t$, see (12). In the marginally recurrent case of coalescing simple random walks in 2D, Lukins, Tribe and Zaboronski LTZ (18) generalized the method of vdBK (00) and proved
$\rho_{N}({\bf x},t)\sim\rho_{1}(t)^{N}\times p_{\rm NC}({\bf x},t),$ (2)
@@
where the correction term $p_{\rm NC}({\bf x},t)$ is the non-collision probability of $N$ independent walks starting from $\bf x$} by time $t$, and it is shown in (CMP, 10, Prop 1.3) that $p_{\rm NC}({\bf x},t)\sim c_{\bf x}(\log t)^{-n(n-1)/2}$. This method can be extended to $d\geq 3$, but fails for $d=1$ due to the strong correlations induced by the recurrence of the underlying random walk. Nevertheless, in the case of 1D coalescing Brownian motions, MRTZ (06) derived the asymptotics for $\rho_{N}({\bf x},t)$, relying crucially on the integrability of the system–most notably, the explicit Karlin-McGregor formula.
@@
In contrast to the nearest-neighbor case, much less is known about the $N$-point function for CRW with long-range jumps. The only existing result in this direction is for 1D coalescing $\alpha$-stable processes with $\alpha>1$, where the density $\rho_{1}(t)$ has asymptotic order $t^{1/\alpha}$, derived as a corollary of (EMS, 13, Thm 6.3 and Prop 2.5) and the scaling property of $\alpha$-stable process.
@@
Motivated by this gap, we study coalescing heavy-tailed random walks on $\mathbb{Z}$}, where the jump increments lie in the domain of attraction of $\alpha$-stable distributions for $0<\alpha\leq 1$. Our results reveal that a form of (2) remains valid in this setting. Notably, $\alpha=1$ is the critical point for recurrence/transience, and our study encompasses a class of recurrent random walks.
@@
Notation. Throughout the paper, we use $c,C$ to denote various positive constants that may even vary within the same line. For simplicity, when summing over an entire set, we omit the explicit domain and only display the summation variable. For example, for ${\bf x}\in{\mathbb{Z}}^{N}$, we write $\sum_{{\bf x}}$ to mean $\sum_{{\bf x}\in{\mathbb{Z}}^{N}}$. For two functions $f(t)$ and $g(t)$, we adopt the convention that
@@
To avoid ambiguities, for $k$-th power of a function $f(t)$, we write $f(t)^{k}$ instead of $f^{k}(t)$. For example, in (2) we do not write $\rho_{1}^{N}(t)$ but $\rho_{1}(t)^{N}$.
@@ 1.1 Model and main result
It is well known Lig (05) that CRW can be constructed by sequentially adding independent random walks, halting each upon its first collision with any of the previously added paths. Precisely, the coalescing heavy-tailed random walks $(\xi_{t})_{t\geq 0}$ on $\mathbb{Z}$} can be defined as a $\{0,1\}^{\mathbb{Z}}$-valued Markov process by specifying its generator thanks to (Lig, 05, Chapter I, Thm 3.9). For $\xi\in\{0,1\}^{\mathbb{Z}}$ and ${\bf x}=(x_{1},\ldots,x_{n})\in{\mathbb{Z}}^{n}$, write $\xi(x_{1},\ldots,x_{n}):=(\xi(x_{1}),\ldots,\xi(x_{n}))\in\{0,1\}^{n}$. We simplify notation by writing $\xi({\bf x})=1$ if $\xi({\bf x})=(1,\ldots,1)$, and $\xi=1$ if $\xi(x)=1$ for all $x\in{\mathbb{Z}}$. Additionally, we represent elements of $\{0,1\}^{n}$ as binary strings, for example, $(0,1)$ is written as $01$. With this notation, the generator $\cal G$} of $(\xi_{t})_{t\geq 0}$ applied on local cylinder functions $f$, i.e., functions that only depend on finitely many coordinates of $\xi$, is given by
$\begin{array}[]{l}\displaystyle{\cal G}f(\xi)=\sum_{x,y\in{\mathbb{Z}}}p(y-x)1_{\{\xi(x,y)=10\}}\big\{f(\xi+e_{y}-e_{x})-f(\xi)\big\}\\[5.0pt] \hskip 62.59596pt\displaystyle+\sum_{x,y\in{\mathbb{Z}}}p(y-x)1_{\{\xi(x,y)=11\}}\big\{f(\xi-e_{x})-f(\xi)\big\},\end{array}$ (3)
where $e_{x}(y)=1_{\{x=y\}}$, and $p(\cdot)$ with $p(0)=0$ and $\sum_{x}p(x)=1$ denotes the jump kernel, which is in the domain of attraction of an $\alpha$-stable distribution. Without loss of generality, we assume that $p(\cdot)$ is aperiodic. By (IL, 71, Thm 2.6.1), it satisfies
$\sum_{|y|\geq x}p(y)=\frac{L(x)}{x^{\alpha}}\big(1+o(1)\big)\quad\text{as}~x\to\infty,$ (4)
@@
To avoid complexity, we assume that $p$ is symmetric throughout the paper. Although, in principle, our results should remain valid in the non-symmetric case, there is a technical obstacle in adapting the present proof, and we therefore leave this extension open (see Subsection 1.3 and Remark 2.8 for further discussion). For the SVF $L$, there exist $1/\alpha$-regularly varying functions $b_{t}:(0,\infty)\to(0,\infty)$ such that $L(b_{t})/b_{t}^{\hskip 1.0pt\alpha}\sim 1/t$. Let $(X_{t})_{t\geq 0}$ be a rate 1 continuous-time random walk with jump kernel $p$ and initial position $X_{0}=0$. It is well known that $X_{t}/b_{t}$ converges weakly to an $\alpha$-stable distribution with probability density function $g_{\alpha}(x)$. By the asymptotic uniqueness of $b_{t}$ (BGT, 89, Thm 1.5.12) and Gnedenko’s local limit theorem (IL, 71, Thm 4.2.1), we can choose a specific $1/\alpha$-regularly varying function
$B_{t}:={\mathbb{P}}(X_{t}=0)^{-1}g_{\alpha}(0)$ (5)
@@
$\frac{L(B_{t})}{B_{t}^{\alpha}}\sim\frac{1}{t}\quad\mbox{and}\quad\sup_{x\in{\mathbb{Z}}}\big|B_{t}{\mathbb{P}}(X_{t}=x)-g_{\alpha}(x/B_{t})\big|=o(1).$ (6)
Recall that $X$ is recurrent if and only if $G(0)=\lim\limits_{x\to\infty}G(0,t)=\infty$, where $G(x,t):=\int_{0}^{t}{\mathbb{P}}(X_{s}=x)\mathrm{d}s$ denotes the Green function of $X$ at $x$ up to time $t$. Following this standard approach, define
$l(t):=\frac{G(0,t)}{g_{\alpha}(0)}=\int_{0}^{t}\frac{1}{B_{s}}\mathrm{d}s,\quad\ell(t):=\frac{B_{t}\hskip 1.0ptl(t)}{t}\quad\mbox{and}\quad l(\infty):=\lim\limits_{t\to\infty}l(t)\in(0,\infty].$ (7)
In the regime $\alpha\in(0,1]$,
$\begin{array}[]{rclcl}\displaystyle X~\text{is (marginally) recurrent}&\Leftrightarrow&l(\infty)=\infty,&&\text{if~}\alpha=1~\text{and}~1/B_{t}~\text{is not integrable,}\\[5.0pt] \displaystyle X~\text{is transient}&\Leftrightarrow&l(\infty)<\infty,&&\text{otherwise.}\end{array}$ (8)
In both cases, $l(t)$ is a SVF (hence we adopt the notation $l$ instead of $G$). Moreover, $\ell(t)$ diverges by Karamata’s theorem and is $(\frac{1}{\alpha}\!-\!1)$-regularly varying. In the marginally recurrent case, we impose an additional technical assumption to ensure that the recurrence is sufficiently weak, in the sense that the tail probability of the first return time decays slowly.
@@
Therefore, we focus on two regimes: either the underlying random walk is transient, or it is marginally recurrent with $B_{t}\in{\cal B}$. In our framework, the condition $B_{t}\in{\cal B}$ cannot be easily removed due to technical reasons; see Subsection 1.3 for further discussion.
[…]
  • If $L(t)\sim(\log t)^{a}$ for some $a<1$, then $B_{t}\sim t(\log t)^{a}$, $l(t)\asymp(\log t)^{1-a}$ and $\ell(t)\asymp\log t$, so that $B_{t}\in{\cal B}$ with any K>1aK>1-a.
  • If $L(t)\sim\exp\{-(\log t)^{\kappa}\}$ for some $\kappa\in(0,1)$, then $B_{t}\sim t\hskip 1.0pt\exp\{-(\log t)^{\kappa}\}$, which yields that $l(t)\asymp(\log t)^{1-\kappa}\exp\{(\log t)^{\kappa}\}$ and $\ell(t)\asymp(\log t)^{1-\kappa}$, and hence $B_{t}\notin{\cal B}$.
Our main result, Theorem 1.3, concerns the asymptotic behavior of the $N$-point correlation function. This relies on two key ingredients as in (2). One is an analysis of the density function $\rho_{1}(t)$ (Theorem 1.1), and the other is an asymptotic result for the non-collision probability of $N$ independent heavy-tailed random walks with fixed starting point $\bf x$} (Theorem 1.2). In fact, we go a step further than (2) by providing sharp estimates for these quantities, establishing explicit error bounds in the results below.
Theorem 1.1 (Density asymptotics)
Let $(\xi_{t})_{t\geq 0}$ be the coalescing heavy-tailed random walks on $\mathbb{Z}$} with initial configuration $\xi_{0}\equiv 1$. Assume the random walk has rate 1 and a symmetric jump kernel $p(\cdot)$ of the form (4) with $0<\alpha\leq 1$. In the marginally recurrent case, further assume $B_{t}\in{\cal B}$ so that $\ell(t)^{K}\geq l(t)$ for some $K$. The density of CRW at time $t$ is
$\rho_{1}(t):={\mathbb{E}}[\xi_{t}(0)]={\mathbb{P}}(\xi_{t}(0)=1).$
Then, there exists some $\eta>0$ such that for any $\varepsilon>0$,
$\rho_{1}(t)=\frac{g_{\alpha}(0)\hskip 1.0ptl(2t)}{t}\big(1+O\big({\cal R}_{1}(t)\big)\big),$ (9)
where the error bound ${\cal R}_{1}(t)$ can be chosen as follows. (1) In the marginally recurrent case where $\alpha=1$, $l(\infty)=\infty$ and $B_{t}\in{\cal B}$, we have ${\cal R}_{1}(t)=l(t)^{-(\frac{1}{2K}-\varepsilon)}$. (2) In the transient case, if $\alpha=1$ and $l(\infty)<\infty$, then ${\cal R}_{1}(t)=\ell(t)^{-1/2}$; and if $0<\alpha<1$, then ${\cal R}_{1}(t)=t^{-\eta}$.
There exists a natural coupling between coalescing random walks and independent random walks so that the two collections of paths are identical up to the first collision time. For convenience, the following theorem is stated for independent walks.
Theorem 1.2 (Non-collision probability asymptotics)
Let ${\bf x}\in{\mathbb{Z}}^{N}$, and let $({\mathbf{X}}_{t})_{t\geq 0}=(X^{1}_{t},\ldots,X^{N}_{t})_{t\geq 0}$, starting from ${\mathbf{X}}_{0}={\bf x}$, be $N$ independent continuous-time random walks with rate $1$ and symmetric jump kernel $p(\cdot)$ of the form (4) with $0<\alpha\leq 1$. In the marginally recurrent case, further assume $B_{t}\in{\cal B}$. Define the first collision time
$\tau_{{\bf x}}:=\inf\{t\geq 0:X^{i}_{t}=X^{j}_{t}~\text{for some}~1\leq i<j\leq N\},$ (10)
and set the non-collision probability $p_{\rm NC}({\bf x},t):={\mathbb{P}}(\tau_{{\bf x}}>t)$. Denote $\upsilon_{N}:={N\choose 2}=\frac{N(N-1)}{2}$. Then, there exists $c_{\bf x}$ depending only on $\bf x$} such that
$p_{\rm NC}({\bf x},t)=\frac{c_{\bf x}}{l(2t)^{\upsilon_{N}}}\Big(1+O({\cal R}_{\rm NC}(t))\Big),$ (11)
where the error bound ${\cal R}_{\rm NC}(t)={\cal R}_{\rm NC}({\bf x},t)$ also depends on $\bf x$}, and can be chosen as follows.
(1) In the case of either $(a)$ marginal recurrence where $\alpha=1$, $l(\infty)=\infty$ and $B_{t}\in{\cal B}$, or $(b)$ marginal transience with $\alpha=1$ and $l(\infty)<\infty$, we can choose
${\cal R}_{\rm NC}(t)=C_{\bf x}\int_{t}^{\infty}\frac{1}{u\hskip 1.0pt\ell(u)^{2-\varepsilon}}\mathrm{d}u=o\big(\ell(t)^{-2+\varepsilon}\big)\quad\text{for any}~\varepsilon>0.$
(2) In the transient case with $0<\alpha<1$, we can choose ${\cal R}_{\rm NC}(t)=C_{\bf x}t^{-\eta}$ for any $\eta>1/\alpha-1$.
We can now present the asymptotic behavior of the $N$-point correlation function of CRW, in line with the spirit of (2).
@@
Let $(\xi_{t})_{t\geq 0}$ be the coalescing heavy-tailed random walks on $\mathbb{Z}$} with initial configuration $\xi_{0}\equiv 1$. Assume the random walk has jump rate 1 and symmetric jump kernel $p(\cdot)$ of the form (4) with $0<\alpha\leq 1$. In the marginally recurrent case with $\alpha=1$ and $l(\infty)=\infty$, further assume that $B_{t}\in{\cal B}$. Denote $\upsilon_{N}={N\choose 2}=\frac{N(N-1)}{2}$. For $N>0$ and ${\bf x}=(x_{1},\ldots,x_{N})\in{\mathbb{Z}}^{N}$, let
$\rho_{N}({\bf x},t):={\mathbb{P}}(\xi_{t}({\bf x})=1)$ (12)
@@
be the $N$-point correlation function. Recall $l(t),\ell(t)$ from (7), and recall $K$ such that $\ell(t)^{K}\geq l(t)$. Recall $\rho_{1}(t)$ and $p_{\rm NC}({\bf x},t)$ from Theorems 1.1 and 1.2, respectively. Then, there exist some $\eta>0$ and constant $c_{\bf x}$ as given in Theorems 1.2 depending on $\bf x$}, such that for any $\varepsilon>0$,
$\rho_{N}({\bf x},t)=\frac{c_{\bf x}\hskip 1.0ptg_{\alpha}^{N}(0)}{t^{N}\hskip 1.0ptl(2t)^{\upsilon_{N}-N}}\big(1+O\big({\cal R}_{N}(t)\big)\big)\sim\rho_{1}(t)^{N}\times p_{\rm NC}({\bf x},t),$ (13)
where the error bound ${\cal R}_{N}(t)={\cal R}_{N}({\bf x},t)$ also depends on $\bf x$}, and can be chosen as follows.
(1) In the marginally recurrent case with $\alpha=1$, $l(\infty)=\infty$ and $B_{t}\in{\cal B}$, for nay $\varepsilon>0$, we can choose ${\cal R}_{N}(t)=C_{\bf x}l(t)^{-(\frac{1}{2K}-\varepsilon)}$;
(2) In the marginally transient case with $\alpha=1$ and $l(\infty)<\infty$, we have ${\cal R}_{N}(t)=C_{\bf x}\ell(t)^{-(\frac{1}{2}-\varepsilon)}$;
(3) In the transient case with $0<\alpha<1$, we have ${\cal R}_{N}(t)=C_{\bf x}t^{-\eta}$ for any $\eta>1/\alpha-1$.
Remark 1.4
The constant $c_{\bf x}$ appearing in Theorems 1.2 and 1.3 depends on both $N$ and the configuration $\bf x$}. It admits an integral representation whose integrand involves the finite-time behavior of the process $({\mathbf{X}}_{t})_{t\geq 0}$, see (85) and the discussion above. However, since this expression is not particularly informative for our purposes, and to keep the statements of the theorems reasonably concise, we omit the explicit formula for $c_{\mathbf{x}}$. We note, nevertheless, that $c_{\mathbf{x}}$ may vanish as $N\to\infty$. Therefore, our results do not apply in the regime where one first sends $N\to\infty$ and then $t\to\infty$.
Remark 1.5
The most interesting case is the recurrent underlying random walk with $\alpha=1$, $l(\infty)=\infty$ and $B_{t}\in{\cal B}$ in Theorems 1.1 and 1.3. This mirrors the 2-dimensional simple CRW studied in BG (80); CP (04); LTZ (18). Indeed, in the 2D setting with our notation one takes $l(t)=\ell(t)=\log t$ and $K=1$, and then $\rho_{1}(t)$ and $\rho_{N}(t)$ exhibit the same leading asymptotics as in Theorems 1.1 and 1.3. Moreover, the error bound in $\rho_{1}(t)$ can be improved from $O\big(l(t)^{-(1/2-\varepsilon)}\big)$ with any $\varepsilon>0$ to $O\big(l(t)^{-1/2}\big)$.
1.2 Proof strategy
We first outline the proof of Theorems 1.1. For high-dimensional ($d\geq 3$) coalescing random walks, van der Berg and Kesten vdBK (00, 02) developed a framework to analyze the asymptotics of the system density $\rho_{1}(t)$ by deriving an ODE through the generator calculations. This robust method has since been applied for CRW on transitive graphs HLYZ (22) and to coalescing systems with multiple particle types TZ (24). Lukins, Tribe, and Zaboronski LTZ (18) adapted this framework to the barely recurrent case in two dimensions. Furthermore, it is adapted by Lukins, Tribe and Zaboronski LTZ (18) to two-dimensional coalescing random walks which is barely recurrent. In our setup, we start by applying the generator (3) to the density function to obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=-\sum_{y}p(y){\mathbb{P}}(\xi_{t}(0,y)=11)$ (14)
To have particles both at sites 0 and $y$ at time $t$, we can trace back the trajectories for their ancestors (which are typically unique) over a time interval $s$ (with $s\ll t$). In effect, we consider these as two backward random walks that have not coalesced between tst-s and $t$. By applying the Markov property at time tst-s, we can decompose
${\mathbb{P}}(\xi_{t}(0,y)=11)\displaystyle\approx\sum_{z_{1},z_{2}}{\mathbb{P}}(\xi_{t-s}(z_{1},z_{2})=11)\hskip 1.0pt{\mathbb{P}}(\widehat{X}^{0,y}_{s}=(z_{1},z_{2}),\hat{\tau}_{y}>s),$ (15)
where $(\widehat{X}^{0,y}_{t})_{t\geq 0}$ denotes two independent backward random walks starting from 0 and $y$, respectively, and $\hat{\tau}_{y}$ is their first collision time. Choosing $s=s(t)$ with $1\ll s\ll t$ appropriately ensures that the backward walks are typically far apart at time $s$ (i.e., $|z_{1}-z_{2}|\asymp B_{s(t)}$), so that ${\mathbb{P}}(\xi_{t-s}(z_{1},z_{2})=11)\approx{\mathbb{P}}(\xi_{t-s}(z_{1})=1)\times{\mathbb{P}}(\xi_{t-s}(z_{2})=1)=\rho_{1}^{2}(t-s)$. Substituting this into (15) and using $\rho_{1}(t-s)\approx\rho_{1}(t)$ (since tst-s is close to $t$) yields
${\mathbb{P}}(\xi_{t}(0,y)=11)\approx\rho_{1}^{2}(t-s){\mathbb{P}}(\hat{\tau}_{y}>s)\approx\rho_{1}^{2}(t){\mathbb{P}}(\hat{\tau}_{y}>t),$ (16)
where we used the fact that the random walk is either transient or barely recurrent (in the sense that $B_{t}\in{\cal B}$) so that ${\mathbb{P}}(\hat{\tau}_{y}>s)\approx{\mathbb{P}}(\hat{\tau}_{y}>t)$ for a suitable choice of $s=s(t)$. Combining (14) and (16) produces a simple ODE for $\rho_{1}(t)$.
Note that (15)-(16) shows an approximation for the two-point function. One can extend these ideas to general ${\bf x}\in{\mathbb{Z}}^{N}$ to prove Theorem 1.3. First, we can establish the asymptotic independence
${\mathbb{P}}(\widehat{X}^{{\bf x}}_{s}={\bf z},\hat{\tau}_{{\bf x}}>s)\approx{\mathbb{P}}(\widehat{X}^{{\bf x}}_{s}={\bf z}){\mathbb{P}}(\hat{\tau}_{{\bf x}}>s),$ (17)
$\rho_{N}({\bf x},t)\approx\sum_{\bf z}\rho_{N}({\bf z},t-s){\mathbb{P}}(\widehat{X}^{{\bf x}}_{s}={\bf z},\hat{\tau}_{{\bf x}}>s)\approx\rho_{1}^{N}(t-s)\hskip 1.0pt{\mathbb{P}}(\hat{\tau}_{{\bf x}}>s)\approx\rho_{1}^{N}(t)\hskip 1.0pt{\mathbb{P}}(\hat{\tau}_{{\bf x}}>t),$ (18)
It remains to establish Theorem 1.2, which concerns the non-collision probability $p_{\rm NC}(t)=p_{\rm NC}({\bf x},t)={\mathbb{P}}(\tau_{{\bf x}}>t)$ for $N$ independent random walks $({\mathbf{X}}_{t})_{t\geq 0}$. The first step is to derive a crude bound Lemma 3.2, so that we can choose a suitable $s=s(t)\ll t$ satisfying $p_{\rm NC}(s)\approx p_{\rm NC}(t)$. Then, we derive an ODE for $p_{\rm NC}(t)$ in a manner analogous to the previous arguments, as also done in LTZ (18) for 2D simple CRW. Since the decrease in $p_{\rm NC}(t)$ is due to collisions between any two random walks $X^{i}$ and $X^{j}$, by the symmetry of $p$ and translation invariance we have